OpenStudy (anonymous):

Find the derivative of f(x)=7 by the LIMIT PROCESS

zepdrix (zepdrix):

\[\large \lim_{h \rightarrow 0}\frac{f(x+h)-f(x)}{h}\] \[\large f(x)=7\]To find \(f(x+h)\), we'll plug in (x+h) everywhere we see an x in the function. Which gives us,\[\large f(x+h)=7\]There are no x's in this function so it doesn't change. :) From there just plug the pieces in! :D